zgilley2:

Which best describes the relationship between the successive terms in the sequence shown? 2.4, –4.8, 9.6, –19.2 The common difference is –7.2. The common difference is –2.4. The common ratio is –2.0. The common ratio is –0.5.

RyeBread1:

cant you look this up online?

darkknight:

just find the common difference between the two

jhonyy9:

2,4*(-2)=-4,8 (-4,8)*(-2)=9,6 9,6 *(-2)= -19,2 hope this help getting easy the correct answer ...